c) Consider John who consumes two goods, (X and Y), with prices š‘ƒš‘„ = š‘$35, š‘ƒš‘¦ = š‘$25 and income I =N$1500 i) Construct budget constraint [3 marks] ii) Draw Johnā€™s budget line with good X on the horizontal axis. [3 marks] iii) Use a graph to show the effect of an increase in income from N$1500 to N$2000. [3 marks] iv) What will happen to the slope of the budget line if the price of good X decreases to N$18? [5 marks]


1
Expert's answer
2021-05-10T16:16:25-0400

i) The budget constraint will be:

"1500=35X+25Y"

ii) The budget line for an income of $1,500 is as follows:



iii)Ā The budget line for an income of $2,000 is as follows:



(iv)

given budget line is

"1500 = 35X +25Y"

TheĀ slopeĀ ofĀ theĀ budgetĀ lineĀ Ā equalsĀ theĀ priceĀ ratioĀ ofĀ twoĀ products.

SlopeĀ ofĀ budgetĀ lineĀ hereĀ "= \frac{Px}{Py}= \frac{35}{25}=\frac{7}{5}



=1.4"

Now when the price of X has changed to $18 and price of Y remains constant.

Therefore slope"=\frac{Px}{Py}=\frac{18}{25}=0.72"

Since slope of budget line is directly related to the price of good X, the fall in price of X will lead to fall in slope of budget line.